how to solve for each of these logs:
4^ln(3x+2) = 13
12^e0.03x = 15^e0.01x

Answer by lwsshak3(11628)

how to solve for each of these logs:
4^ln(3x+2) = 13
12^e0.03x = 15^e0.01x
**
4^ln(3x+2) = 13
take log of both sides
ln(3x+2)ln4=ln13
ln(3x+2)=ln13/ln4=1.8502
convert to exponential form: base(e) raised to log of number(1.8502)=number(3x+2)
e^1.8502=3x+2=6.3611
3x=4.3611
x=4.3611/3=2.1201
..
12^e0.03x = 15^e0.01x
take log of both sides
e0.03xln12=e0.01xln15
e0.03x/e0.01x=ln15/ln12=1.0898
3≠1.0898
no solution